3. Suppose that 88% of college men and 84% of college women were employed last summer, based on a sample survey that interviewed SRSs of 400 college men and 400 college women. The two samples were independent.

a) What is the approximate distribution of the proportion pF of women who worked last summer? What is the approximate distribution of the proportion pM of men who worked? Write your answer in a Text Box.

b) The survey wants to compare men and women. What is the approximate distribution of the difference in the proportions who worked, pM - pF? Explain and write your answer in a Text Box.

c) What is the probability that in the sample a higher proportion of women than men worked last summer?
Answer #1

Solution:

An Swea: Given Dala is wonen ten ni - H00 ng - 400 Bi - 0.88 B, 0.54 (a) proposition foor women P2 = 0.84 Standard error - Puно-о – ? 2) а = О - О 24 = P (24 - 1. 632 ) – о. 5 - P ( - (- 632 4 20) - б. 5 - о. н- 8- - б: 0516.